# UINav: A maker of UI automation agents

### v1

@@ 9 Limitations
Figure 12: An agent selects an action to turn off notification dot (a) when the switch is on, and (b) when the switch is already off. The texts in red (click in a) and (wait in b)) are the actions selected by the agent.
10 Conclusions
In this paper, we presented a demonstration-based system for building UI automation agents capable of performing many tasks in real-world applications. The process we have proposed requires small human effort and no coding skills, while the resulting agents achieve near perfect success rate on previously-seen tasks and also generalize well to new tasks and applications. The trained neural networks are small and fast, thus being suitable for mobile devices.
Acknowledgements
We would like to thank Chris Rawles, Alice Li, Yifan Chang, Marybeth Fair, and Nghi Doan for the help on testing the system.
Appendix A Appendix
A.1 Case study of agent capabilities
A.1.1 Sending an email with multiple text inputs
@@
Figure 10 shows the image sequence of a UINav agent completing the “send email” task. The task utterance is “send an email to uinav@gmail.com with the subject: Events for the Week and the content: Hi, you are invited to attend our weekly meeting from 4:00pm to 5:00pm on Tuesday (EST - New York)”.
A.1.2 Sending a message to the correct recipient
@@
Figure 11 compares two cases of an agent sending messages. The images are deliberately modified to hide the real names of the recipients. Both (a) and (b) are in the message view of the app but of different recipients, Tom in (a) and Jerry in (b), while the utterance specifies the recipient to be Jerry. The agent correctly recognizes the difference and selects the correct action for both cases: pressing the back button at the top left for (a) and typing the content of the message at the bottom for (b). Note that it is the title bar that contains the information on the current recipient. We believe that it is due to the self-attention of the Transformer encoder that the agent learns whether the text of the title bar matches the recipient is a critical signal in these states.
A.1.3 Understanding the relationship between text label and switch
@@
Figure 12 shows how the UINav agent selects actions to turn off notification dot in two cases: (a) when the switch is on and the agent selects the action to click the text label of "Notification dot on app icon", and (b) when the switch is already off and the agent chooses to wait for the referee to terminate the task. Note that the text label of "Notification dot on app icon" and its switch are independent UI elements in the screen representation, and there are multiple switches on the screen with identical attributes except for their positions and states. The agent learns their relationship probably by the relative positions (horizontally aligned).

### v2

@@ Appendix A Appendix
Figure 6: UINav agent searches in YouTube. The pink arrows highlight the agent’s actions that are also annotated by red boxes and texts. To start using the search bar the agent must first dismiss popups (twice) and clear the search bar. (a) Clicks the back button to dismiss a popup ads; (b) Clicks "X" to dismiss the install page of Audible; (c) Clicks "X" to erase the previously entered search phrase “something”; (d) Focuses on the search bar to enter a new search term.
A.1 An example task: search in YouTube
Fig. 6 shows the UINav agent searching in YouTube. The agent dismisses popups twice (a) and (b) to reveal the search bar. It then clicks the "X" button to erase the previous search phrase “something” (c). The system does not reach the desired start state for a search until the screen shown in (d), where the agent sets the focus on the search bar to then enter the search term.
Fig. 4 shows the SEARCH task required over 1700 task demonstrations because it must work for 100 or more different apps and websites. All other tasks are specific to a single app and thus require fewer samples, 33 on average.
A.2 Action space
Table 3: UINav action space.
Element click <elem> Clicks the center of the specified element.
actions focus_and_type <elem,text> Sets focus on the specified element, types the specified text, and optionally presses Enter.
dismiss <elem> Clicks outside of the specified element.
Global wait Waits until the next observation is received.
actions back Goes back to the previous app screen.
scroll <left|right|up|down> Scrolls in the specified direction.
open_app <app_name> Launches the specified application.
A.3 Referee model
The referee is a recurrent neural network (RNN)-based model (Fig. 7). The attention over Transformer-encoded UI elements is similar to that of the agent model, except that the query is the input utterance concatenated with the action history (the action performed in the previous step and its outcome). Although action history could be derived from previous screen representations, feeding it as input directly makes it less challenging as the referee does not have to learn it. The output of the attention module is then fed into a gated recurrent unit (GRU) Cho et al. (2014). The GRU takes this along with the previous internal hidden state as inputs to predict the current status of the step: (1) SUCCESSFUL: the task is completed and it is successful; (2) FAILED: the task has failed or has reached the maximum number of allowed steps; (3) PENDING: the task is ongoing; or (4) INFEASIBLE: the task cannot be executed (e.g., the task may not be well defined). Failed executions are saved along with all their parameters and passed to the demonstrator.
Figure 7: The architecture of the UINav referee model.
A.4 Macro actions
In UINav, every action is executed as a small program that is composed of lower level operations with status checks. Such a program is referred to as macro. Macro actions abstract the system-specific details, thus making it possible to build cross-platform agents and simplifying the agent’s logic. Each macro action is implemented following a state transition graph. Fig. 8 shows the state transition graph for most macro actions that result in screen changes, such as click and back. It starts at S0, and transitions among the other states according to incoming events, such as Action dispatched and Screen changed, and exits either successfully (S6) or with a failure (S5). The graphs of other macro actions are similar.
Another advantage of using macro actions is that they package highly dependent, low-level actions. Fig. 9 shows an example. The focus_and_type action (inspired from MiniWoB Shi et al. (2017)) consists of 4 low-level actions: clicking the input field to obtain focus, waiting for the blinking cursor to appear, typing the text in the field, and (optionally) pressing Enter. (Note that large arrows in purple are drawn to highlight interesting areas.)
A.5 Implementation
We built UINav for the Android platform. However, our design is applicable to other platforms and some of our techniques (e.g., macro-actions and screen representation) are specifically designed to be platform agnostic. Both the agent and the referee models are implemented in TensorFlow. We employ two inference modes, off-device and on-device. During development we use the Python API of TensorFlow to test the models off-device. Once stable, the models are converted to TensorFlow Lite (tflite) for on-device inference. Both agent and referee models utilize the same pre-trained language model to encode utterances and texts appearing on screens. We choose the smallest model, L-2_H-128_A-2, of SmallBERT Turc et al. (2019), and convert it to a 17.6MB tflite model. Note that no quantization is applied during the tflite conversion of any of the above models. For efficiency, the sentence encoding computation of the agent and referee models are shared.
A.6 UINav Console
To collect demonstrations, we have developed a dedicated application, the UINav Console, that can be seen in the right-half side of the screenshots in Fig. 1012. At each step of a demonstration, a user specifies a macro action, including action type, referenced element, and action argument (if any), and then requests execution of the action.
A.7 Model training details
Training the agent model.
For the agent model, demo augmentation happens dynamically with a 1% probability for a sample to remain unchanged. The model is optimized by an Adam optimizer with a fixed learning rate of 1e-3. Initially a training runs up to 100,000 samples and can be terminated earlier if the test accuracy stabilizes. If new demonstrations are added, the agent will be trained with additional 20,000 samples. It is trained on CPU or GPU with a batch size of 256.
Training the referee model.
For the referee model, each demonstration is augmented to 10 samples at a pre-possessing stage. The model is optimized by an Adam optimizer with a fixed learning rate of 1e-3. A training takes up to 30 epochs and can be terminated earlier if the test accuracy stabilizes. It is trained on CPU or GPU with a batch size of 128.
A.8 Case study of agent capabilities
@@
Figure 12: An agent selects an action to turn off notification dot (a) when the switch is on, and (b) when the switch is already off. The texts in red (click in a) and wait in b)) are the actions selected by the agent.
Sending an email with multiple text inputs.
@@
Fig. 10 shows the image sequence of a UINav agent completing the “send email” task. The task utterance is “send an email to uinav@gmail.com with the subject: Events for the Week and the content: Hi, you are invited to attend our weekly meeting from 4:00pm to 5:00pm on Tuesday (EST - New York)”.
Sending a message to the correct recipient.
@@
Fig. 11 compares two cases of an agent sending messages. The images are deliberately modified to hide the real names of the recipients. Both (a) and (b) are in the message view of the app but of different recipients, Tom in (a) and Jerry in (b), while the utterance specifies the recipient to be Jerry. The agent correctly recognizes the difference and selects the correct action for both cases: pressing the back button at the top left for (a) and typing the content of the message at the bottom for (b). Note that it is the title bar that contains the information on the current recipient. We believe that it is due to the self-attention of the Transformer encoder that the agent learns whether the text of the title bar matches the recipient is a critical signal in these states.
Understanding the relationship between text label and switch.
@@
Fig. 12 shows how the UINav agent selects actions to turn off notification dot in two cases: (a) when the switch is on and the agent selects the action to click the text label of "Notification dot on app icon", and (b) when the switch is already off and the agent chooses to wait for the referee to terminate the task. Note that the text label of "Notification dot on app icon" and its switch are independent UI elements in the screen representation, and there are multiple switches on the screen with identical attributes except for their positions and states. The agent learns their relationship probably by the relative positions (horizontally aligned).
A.9 Apps and websites used in data collection
The full list of Android apps and websites that are used in our data collection is as follows:
Facebook Messenger, TikTok, Instagram, WhatsApp, Amazon Shopping, Facebook, Walmart, Spotify, Pandora, Amazon Prime Video, Google Play Games, Wish, Pinterest, Google Messages, Target, Poshmark, Waze, Twitter, Wayfair, google.com, Google Play Store, Seamless, YouTube, Reddit, Ebay, Etsy, Soundcloud, Tasty, Gmail, Contacts, Android Auto, YouTube Music, Snapchat, Tubi TV, Shop, News Break, Cash App, Pluto TV, Uber, Burger King, Roku, Amazon Alexa, Life 360, HBONow, ESPN, iHeartRadio, Nike, Amazon Photos, Letgo, Walmart Grocery, Weather App, Google News, Files, Home Screen, Google Docs, DoorDash, Google Photos, AirBnB, AliExpress, Amazon Music, Apple Music, Audible, Chewy, Chik Fil A, Costco, Dollar General, Google Drive, Dunkin Donuts, Google Earth, Emoji Home, Family Dollar, wikipedia on firefox, Food Network, GroupMe, Groupon, GrubHub, Instacart, KeepNotes, King James Version, Kroger, Likee, LinkedIn, fb Lite, Lyft, Maps, OfferUp, Phone, Pixaloop, Scanner, SHEIN, Skype, SmartNews, Starbucks, thredUp, Ticket Master, Walgreen’s, Yahoo Mail, Yelp, YouTube Kids, Zedge, Zelle, Zillow, wikipedia.org, youtube.com, yahoo.com, facebook.com, live.com, reddit.com, bing.com, linkedin.com, Sam’s Club, discord, GoodRx, Outlook, Breaking US News, Lucky Go, CNN, Postmates, Transit, Sephora, target.com, twitter.com, irs.gov, craigslist.org, homedepot.com, Recipes Home, Zillow, and Dialer.